OpenStudy (anonymous):

The tangent line to the graph of g(x) = x^3 at the point (x, y) has slope = 3x^2. g (3) =? The slope of the curve at x = 3 is? The equation of the tangent line to the graph of g(x) = x^3 at x = 3 is y =?

myininaya (myininaya):

g(x)=x^3 g'(x)=3x^2 <-slope=derivative (whatever!) g'(3) means slope at x=3 g'(3)=3(3)^2 <- slope at x=3 (we will call this m) ----- lines have form y=mx+b we know a point on this line (x,y)=(3,3^3) plug m,x,y in to find b

OpenStudy (anonymous):

thank you for the help!